### v1

# Efficient generation of divergent and collimated hot electrons via a novel multi-beam two-plasmon decay and stimulated Raman scattering mechanism

Abstract
In inertial confinement fusion (ICF) implosions, the preheating risks associated with hot electrons generated by laser plasma instabilities (LPI) are contingent upon the angular characteristics of these hot electrons for a given total energy. Using particle-in-cell simulations, we reveal a novel multi-beam collaborative mechanism of two-plasmon decay (TPD) and stimulated Raman scattering (SRS), and investigate the angular variations of hot electrons generated from this TPD-SRS instability (TSI) driven collectively by dual laser beams with varying incident angles $\theta_{in}$ ($24^{\circ}$ to $55^{\circ}$ at the incident plane) for typical ICF conditions. In the simulations with $\theta_{in}\gtrsim 44^{\circ}$, TSI emerges as the dominant mechanism responsible for hot electron generation, leading to a wide angular distribution of hot electrons that exhibit both pronounced divergent and collimated components. The common Langmuir wave associated with TSI plays a crucial role in accelerating both components. By properly modeling the TSI common wave gains, we establish scaling relations between these gains and the energies of collimated and divergent hot electrons. These relations reveal that the divergent hot electrons are more sensitive to variations in gain compared to the collimated electrons. Additionally, the calculated gains qualitatively predict the asymmetry in hot electron angular distributions when the density gradients deviate from the bisector of the laser beams. Our findings offers insights for hot electron generation with multiple beams, potentially complementing previous experiments that underscore the critical role of overlapped intensity from symmetric beams within the same cone and the dominance of dual-beam coupling.
preprint: AIP/123-QED
I Introduction
In inertial confinement fusion (ICF), a spherical capsule is imploded by multiple high-power laser beams to compress the central fuels to extreme density and pressure, thereby initiating self-sustained fusion reactionsAtzeni and Meyer-ter Vehn (2004). During the implosion, hot electrons (with energy above $\sim 50$ keV) produced by laser plasma instabilities(LPI)W.L.Kruer (2003) can preheat the uncompressed fuel and degrade the implosionChristopherson et al. (2021); Solodov et al. (2022). In most ICF schemesNuckolls et al. (1972); He et al. (2016); Zhang et al. (2020a); Lan (2022), such hot electrons need to be reduced to improve the implosion, except for shock ignitionBetti et al. (2007) (SI), in which hot electrons generated during the spike pulse can enhance the shock pressureBaton et al. (2020); Tentori et al. (2021) and enhance the compression. While ignition has been achieved using the indirect-drive scheme on the NIF laser facilityAbu-shawareb (2024) with hot electron threats minimized by unique experimental configurations, the pursuit of higher energy gainsBatani et al. (2023); Sui and Lan (2024) and alternative ignition schemesGopalaswamy et al. (2024); Williams et al. (2024); He et al. (2016); Zhang et al. (2020a); Lan (2022); Betti et al. (2007) still demands effective control and thorough characterization of hot electron effects. Thus, reliably characterizing the properties of hot electrons is crucial to evaluate their potential detrimental or beneficial impacts.
Previous ICF experiments have explored the hot electron energy and temperature scaling relations under various laser-plasma conditionsMichel et al. (2013); Froula et al. (2012). One primary source of these hot electrons is two-plasmon decay (TPD), a major LPI in both directSeka et al. (2009); Turnbull et al. (2020) and indirectRegan et al. (2010) ICF schemes. In TPD, one laser wave decays into two electron plasma waves (EPW), efficiently transferring energy to the EPWsW.L.Kruer (2003), which can accelerate background electrons. In addition to experimental studies, scaling laws have been examined through first-principle particle-in-cell (PIC) simulationsCao et al. (2022); Cao and Ren (2023), considering both TPD and stimulated Raman scattering (SRS)—another key LPI that generates hot electronsRosenberg et al. (2018); Dewald et al. (2016). In SRS, a laser wave decays into an EPW and an electromagnetic wave (EMW)W.L.Kruer (2003).
Various approaches have been developed to mitigate hot electron generation, including advancements in target designFollett et al. (2016) and broadening of laser bandwidthsYao et al. (2024); Lei et al. (2024); Wang et al. (2024); Liu et al. (2024); Ma et al. (2021); Follett et al. (2021); Gao et al. (2020); Zhao et al. (2017). Nevertheless, uncertainties persist regarding the angular distribution of hot electrons. Given hot electron energy and temperature, the impact of hot electrons depends on their angular characteristics. In conventional center hot-spot ignition for both directChristopherson et al. (2021) and indirect driveDewald et al. (2016) ICF, directional hot electrons that propagate towards the capsule center pose a greater preheating threat than those moving laterally. Similarly, the SI scheme benefits from directional hot electrons to augment the shock pressure. Conversely, in the emerging double-cone ignition (DCI) schemeZhang et al. (2020a), hot electrons generated in the compression cone and moving laterally may preheat the ignition cone, forming local pre-plasmas that reduce the energy coupling efficiency of the picosecond ignition laser pulse to the compressed fuelMacPhee et al. (2010); Li et al. (2013).
Experiments have demonstrated that the angular distribution of hot electrons varies with laser illumination geometry and plasma conditions. Early studiesEbrahim et al. (1980) involving single-beam $CO_{2}$ laser pulses interacting with long-scale-length plasmas have revealed that hot electrons are concentrated at $45^{\circ}$ relative to the laser incident direction in the polarization plane. In contrast, more recent experimentsYaakobi et al. (2013) on the OMEGA laser facility have observed a wide divergence of hot electrons during implosions of spherical targets illuminated by 60 laser beams. In subsequent SI-relevant experimentsZhang et al. (2020b) on the OMEGA-EP laser facility, where intense single-beam laser pulses (with intensities of $5-10\times 10^{15}W/cm^{2}$) interacted with planar targets, the measured hot electrons associated with both TPD and SRSLi et al. (2020) exhibited directional characteristics. A more comprehensive experimental studyLiu et al. (2017) using single laser beam at varying intensities has shown that the angular distributions of hot electrons varies depending on the specific laser and target configurations. The broad angular spread of hot electrons is generally attributed to the wide EPW spectrum in the nonlinear phase of TPDMyatt et al. (2014). Nevertheless, the mechanisms behind the generation of hot electrons with various angular features remain insufficiently understood. A key factor influencing these angular characteristics is the laser beam geometry, as the daughter EPWs in TPD and SRS are primarily driven by wave vector matching across multiple laser beams in current ICF experimentsMyatt et al. (2014); Michel et al. (2015).
In this article, we investigate the angular variations of TPD-generated hot electrons in relation to laser illumination geometry using fully kinetic PIC simulations and theoretical analysis. Our results show that laser geometry is the primary determinant of the angular features of hot electrons. In our study, we specifically focus on dual laser beams symmetric to the density gradient, while typical cases with asymmetric incidence are preliminarily analyzed using the theoretical model developed for symmetric configurations, with adjustments to account for asymmetry. There are two key reasons for emphasizing the symmetric dual beam cases. First, experiments have demonstrated that the overlapped intensity of laser beams in the same incident cone plays a dominant role in hot electron generationMichel et al. (2013); Regan et al. (2010). Our study using dual laser beams with identical incident angles is extendable to more beams within the same cone due to angular symmetry. Second, even though the total overlapped intensity in a cone matters most in hot electron generation, typical OMEGA-EP experiments have shown the dominance of dual-beam coupling rather than more beams in TPD growthMichel et al. (2012). We study incident angle $\theta_{in}$ ranging from $24^{\circ}$ to $55^{\circ}$, as $\theta_{in}\sim 60^{\circ}$ brings the laser reflection density near the quarter-critical density $n_{c}/4$, lowering the TPD threshold--- -a scenario already addressed in previous workLian et al. (2022); Zhou et al. (2023). Higher $\theta_{in}$ values shift the reflection density away from $n_{c}/4$ to lower density region, making TPD instability unlikely.
This article is organized as follows. Section II describes the simulation setup. In Section III, we provide simulation results with theoretical explanations and discuss the underlying physics. Section IV offers further discussion on the applicability of our results to realistic experimental conditions. Section V presents the conclusions of this study.
II Simulation setup
Figure 1: (a) Schematic setup used in 2D PIC simulations. Note that the figure is not drawn to scale. (b)The density profile along x-direction.
Table 1: The simulation parameter sets
Index $I_{ovr}$(×$\times 10^{14}W/cm^{2}$) $\theta_{in}(^{\circ})$ $L_{n}(\mu m)$ $T_{e}(keV)$ $T_{i}(keV)$ $\eta$
i 5.0 24,32,40, 150 3 1.5 1.04
44,48,52,
55
ii 3.6 24,40,52 150 2 1 1.12
iii 10.0 24,40,52 75 3 1.5 1.04
iv 2.5 24,40,52 300 3 1.5 1.04
v 5.0 52 150 2 1 1.57
vi 6.3 52 150 2 1 1.97
[…]
The overlapped laser intensity $I_{ovr}$ (twice of the single beam intensity $I_{sig}$), incident angle $\theta_{in}$, initial electron density $n_{e}(x)$, density scale length $L_{n}=n_{e}/(dn_{e}/dx)$ at $n_{e}=n_{c}/4$, as well as electron and ion temperatures $T_{e}$, $T_{i}$ are chosen within specific ranges of (2.5, 10) ×$\times 10^{14}W/cm^{2}$, $(24^{\circ},55^{\circ})$, $(0.2,0.272)n_{c}$, $(75,300)\mu m$, $(2,3)keV$ and $(1,1.5)keV$, respectively, as listed in table (1). We define $I_{sig}=I_{ovr}/2$ as the beam intensity averaged along transverse direction (y-direction) within the e-folding width of the wave field with Gaussian transverse profiles. These conditions are typical for direct-drive ICF experiments with the TPD thresholdSimon et al. (1983) factor $\eta=I_{ovr}L_{n}\lambda/T_{e}/81.86$ from 1.04 (right above threshold $\eta=1$) to 1.97, where $I_{ovr}$, $L_{n}$, $\lambda$ and $T_{e}$ should be in unit of $10^{14}W/cm^{2}$, $\mu m$, $\mu m$ and $keV$, respectively. We deliberately avoid higher $\eta$, or longer $L_{n}$ and higher $T_{e}$ to focus on a regime dominated by multi-beam effects rather than single-beam processes. In our simulations, the reflectivity caused by SRS and stimulated Brillouin scattering (SBS) remains below 3% and 7%, respectively, except for the SBS reflectivity of 11% for the simulation (iii) of table 1 with $\theta_{in}=52^{\circ}$.
At time $t=0$, the p-polarized laser pulses are launched at the left boundary ($x=0$) of the simulation domain. For the simulations with $L_{n}=150\mu m$, the domain size is $L_{x}\times L_{y}=819.2c/\omega_{0}\times 3276.8c/\omega_{0}$ (or $43.5\mu m\times 173.8\mu m$) consisting of $4096\times 16384$ square cells with size $dx=dy=0.2c/\omega_{0}$ ($0.011\mu m$). Each cell is initially distributed 200 numerical particles (100 electrons, 50 fully ionized Carbon ions and 50 Hydrogen ions). It is worth noting that the cell size remains the same for all the simulations with varying domain sizes. The simulations progress with a time step of $dt=0.1414\omega_{0}^{-1}=0.025fs$ until $\sim 6ps$ when TPD has mostly reached a quasi-steady state. We employ open field boundaries and thermal particle boundaries along both x- and y-directions. During the simulations, we record all the hot electrons crossing the boundaries with energies exceeding 50 keV. The electromagnetic fields are diagnosed using the time Fourier transform (TFFT) moduleWen et al. (2019); Cao et al. (2022) with time windows of $\sim 1ps$. This module records the time Fourier-transformed fields throughout the entire space but only within specific frequency ranges: (0.45,0.55)$\omega_{0}$ for Langmuir waves, (0.95,1.05)$\omega_{0}$ for light waves. With frequency filtering, this method provides cleaner signals than the conventional spatial Fourier transform of the “snapshots" and can distinguish the directions of different wave modes. Of course, we still employ conventional “snapshots" diagnostics of all the particles and fields.
III Results and discussions
III.1 Two major characteristics of TPD growth for varying $\theta_{in}$
To illustrate the effects of $\theta_{in}$ on TPD growth, we first present the results of the simulations with identical physical parameters but varying $\theta_{in}$, as listed in table 1(i). These conditions correspond to conventional TPD thresholdSimon et al. (1983) factor $\eta=1.04$, slightly above the threshold of $\eta=1$ to represent a typical TPD-dominant case. As $\theta_{in}$ increases from $24^{\circ}$ to $55^{\circ}$, the TPD growth primarily exhibits two distinct characteristics for relatively smaller ($\theta_{in}<40^{\circ}$) and larger ($\theta_{in}>44^{\circ}$) incident angles, respectively. The two characteristics can be effectively demonstrated by typical cases of $\theta_{in}=24^{\circ}$ and $52^{\circ}$. For these cases, the space (x-direction) and time evolution of electron density perturbations, mainly produced mainly by the TPD daughter EPWs, are plotted in figure (2a) and (2(b), respectively. The color scale in the two figures represents $<\Delta n_{e}>(x,t)$, which is the perturbation level at certain location $x$ and time $t$:
$<\Delta n_{e}>(x,t)=\sqrt{\sum_{i_{y}=1}^{N_{y}}[n_{e}(x,dy\cdot i_{y},t)-n_{e}(x,dy\cdot i_{y},0)]^{2}/N_{y}}$ (1)
where $i_{y}$ is the cell index along $y$-direction, and $N_{y}$ is total the cell number along $y$-direction of the simulation domain.
Figure 2: The space (x-direction) and time evolution of the electron density perturbation $<\Delta n_{e}>$ for the incident angle $24^{\circ}$ and $52^{\circ}$ in simulation series (i) [Table 1]. The $x-$locations of $0.21n_{c}$, $0.23n_{c}$ and $n_{c}/4$ are marked by vertical dashed lines in red, blue and black colors, respectively.
The growth of TPD modes at different times can be demonstrated by $<\Delta n_{e}>(x,t)$ in figure (2). For $\theta_{in}=24^{\circ}$, TPD modes first grow near the $n_{c}/4$ density, then diffuse to $\sim 0.23n_{c}$ and reach the quasi-steady state after $\sim 3ps$. By contrast, TPD modes for $\theta_{in}=52^{\circ}$ initially grow in a much broader space range from $0.22n_{c}$ to $n_{c}/4$, and this range remains until TPD reaches a quasi-steady state.
The different behaviors of $<\Delta n_{e}>(x,t)$ correspond to distinct EPWs of TPD daughter waves with frequency $\sim\omega_{0}/2$. The space spectra of the electric fields with frequency from $0.45\omega_{0}$ to $0.55\omega_{0}$ are plotted in figure 3 for two time intervals: $1$ to $2ps$ and $5$ to $6ps$, representing the linear and nonlinear stages, respectively. To equally reflect the electric fields along $x-$ and $y-$ directions, we define $|E|$ as the magnitude of electric fields in $k_{x}$ and $k_{y}$ phase space with frequency from $0.45\omega_{0}$ to $0.55\omega_{0}$:
$|E|=|E(k_{x},k_{y})|=\sqrt{E_{x}(k_{x},k_{y})^{2}+E_{y}(k_{x},k_{y})^{2}}$ (2)
where $E_{x}$ and $E_{y}$ are the electric fields along $x-$ and $y-$ directions. For $\theta_{in}=24^{\circ}$, TPD grows significantly near the crossing point of the two TPD hyperbolas of the two incident beams in the linear stage [figure (3a)], agreeing with the typical growth mechanism of sharing a forward-going common waveZhang et al. (2014). In the later nonlinear stage, when the quasi-steady state is formed, the spectrum broadens due to nonlinear saturationsMyatt et al. (2012), though the common wave signals remain dominant.
The excitation of such common wave relies on the crossing of the hyperbolas in the forward direction (positive $x-$direction) [figure (3a)(3c)] , which is not satisfied for $\theta_{in}=52^{\circ}$ [figure (3b)(3d)]. In this case, the hyperbolas intersect at the sides with EPWs observed near the intersections. One might intuitively guess they are common waves co-driven by two laser beams via TPD. However, this cannot be the case, as the common wave cannot simultaneously act as the daughter wave with relatively longer $\vec{k}$ (common wave frequency $\omega_{co}>\omega_{0}/2$) for one laser beam and shorter $\vec{k}$ ($\omega_{co}<\omega_{0}$/2) for the other beam when both beams have the same frequency $\omega_{0}$.
We also observe electric field modes at $k_{x}\sim 0$ with discrete peaks at different $k_{y}$’s. The modes with $k_{x}\sim 0$ and $k_{y}\sim 1.6k_{0}$ near the hyperbola have much greater $|k|$ than all the modes for $\theta_{in}=24^{\circ}$. According to the TPD theory, these modes should locate at much lower density than the modes for $\theta_{in}=24^{\circ}$, agreeing with the results in figure (2). Moreover, the overall electric field spectra for $\theta_{in}=52^{\circ}$ are dominated by separated peaks rather than broad signal islands for $\theta_{in}=24^{\circ}$, even in the nonlinear stage. This causes multi-peak angular distributions of hot electrons, which are presented in section III.3.
Figure 3: The space spectra of $|E(k_{x},k_{y})|$ defined by equation (2) for the simulation (i) [table 1] during the time of 1-2ps (a)(b) and 5-6ps(c)(d), and with $\theta_{in}=24^{\circ}$ (a)(c) and $52^{\circ}$ (b)(d). The TPD hyperbola rotated by ±$\pm\theta_{in}$ are plotted as dashed curves, representing the locations corresponding to maximum TPD growth rates for the two incident laser beams. In (a)(c), the colored arrows mark the wave vectors of the laser beams (green and red), common EPW (black) and the other TPD modes (white). The corresponding wave vectors for (b)(d) are given in figure 4(b).
III.2 The excitation mechanism of side TPD modes for relatively larger $\theta_{in}$
For the two characteristics described in the previous section, the one corresponds to $\theta_{in}=24^{\circ}$ is driven through a forward-going common wave and has been well understoodZhang et al. (2014). Thus, we focus on the underlying physics of the side modes excitation for $\theta_{in}=52^{\circ}$. We find that the side EPWs are still common waves, but they are shared by TPD driven by one beam and SRS by the other, instead of two TPDs, so we name it TPD-SRS instability (TSI) in this article. This TSI mechanism can explain most major modes in figure (3b)(3d), especially those side modes with $k_{x}\sim 0$.
To better illustrate the TSI mechanism, we first recall the basic TPD and SRS theories of wave coupling and maximum growth conditions. As typical three-wave coupling processes, both TPD and SRS require matching conditions of frequency $\omega_{0,1,2}$ and wave vector $\vec{k}_{0,1,2}$ to be satisfied for the instability to grow:
$\vec{k_{0}}$ $=\vec{k_{1}}+\vec{k_{2}}$ (3)
$\omega_{0}$ $=\omega_{1}+\omega_{2}$ (4)
where the subscripts 0, 1 and 2 refer to the laser wave and two daughter waves, respectively. The $\omega_{0,1,2}$ and $\vec{k}_{0,1,2}$ must follow the dispersion relations of electromagnetic waves (EMWs) [equation (5)] or EPWs [equation (6)], depending on the wave forms:
$\omega^{2}$ $=\omega_{pe}^{2}+k^{2}c^{2}$ (5)
$\omega^{2}$ $=\omega_{pe}^{2}+3k^{2}v_{te}^{2},$ (6)
where $\omega_{pe}$ is the plasma frequency depending on the density $n_{e}$, $v_{te}=\sqrt{T_{e}/m_{e}}$ denotes the electron thermal speed, where $m_{e}$ is the electron mass.
Given $\omega_{pe}$ and $v_{te}$ (or $n_{e}$ and $T_{e}$), the equations of dispersion relations and matching conditions determine two matching ellipses of TPD and SRS [figure (4a)]. The vectors starting from the origin ending on the ellipses represent all the possible EPW daughter waves satisfying the matching equation (3) to (6). The vectors from the end of $k_{0}$ (the horizontal green arrow) to the blue ellipse mark the corresponding wave vectors of scattered lights via SRS.
Figure 4: (a) Schematic view of the matching ellipses of TPD (black) and SRS (blue) in $k_{x}$ and $k_{y}$ phase space for a laser beam with wave number (green) $k_{0}\sim 0.87\omega_{0}/c$ for a certain $n_{e}$ The red hyperbola $k_{y}^{2}=k_{x}(k_{x}-k_{0})$ corresponds to the wave numbers with the greatest TPD growth rates for different $n_{e}$. For relatively higher $n_{e}$, the ellipses are smaller. (b) The space spectra of $|E(k_{x},k_{y})|$ defined by equation (2) for the simulation (i) [table 1] with $\theta_{in}=52^{\circ}$ during the simulation time of 5-6ps.
With the above theory of matching ellipses, we can conveniently identify all the common waves that can be co-driven by multiple incident laser beams. In figure (4b), We redraw the $|E|$ spectrum for $\theta_{in}=52^{\circ}$ case during 5-6 ps [figure (3d)] with the TPD ellipse (purple dashed line) for the laser beam pointing down-right and the SRS ellipse (brown dashed line) for the laser beam pointing up-right for $n_{e}=0.217n_{c}$ and $T_{e}=3keV$ [table 1(i)]. The two ellipses intersect in $k$ space, suggesting the EPW pointing from the origin to the intersection can be co-driven by the two laser beams via TPD and SRS.
Nonetheless, the formation of such common waves does not guarantee strong growth, which requires high gain of the driven modes. Actually, the matching of the common waves broadly exists for a considerable density range between $0.21n_{c}$ and $0.23n_{c}$, wherever the two matching ellipses intersect. However, only one dominant common wave among these matching modes can grow significantly, as shown in the simulation results [figure (4b)].
To determine the relative significance of all the matching modes, we adopt the Rosenbluth gainRosenbluth, White, and Liu (1973) $G_{R}$, which physically means the amplification factor of a delta-function seed in nonuniform plasmas:
$G_{R}=exp[\frac{\pi\gamma_{0}^{2}}{\kappa^{\prime}V_{1}V_{2}}]$ (7)
where $\gamma_{0}$ is the temporal growth rate of the instability in a homogeneous plasma, $V_{1}$ and $V_{2}$ are the group velocities of the two daughter waves, and $\kappa^{\prime}=\frac{d}{dx}(k_{0x}-k_{1x}-k_{2x})$ denotes the $k$ mismatch factor along the $x-$direction (along the density gradient). The above equation (7) applies to both TPD and SRS with all the variables varying for different $n_{e}$ and $T_{e}$. For simplicity, we do not write this equation separately for TPD and SRS, and all the following equations apply for both TPD and SRS unless otherwise stated.
To provide clarity, we point out that the equation (7) of Rosenbluth gain was developed with WKB approximationsRosenbluth, White, and Liu (1973), which basically requires that all the involved EPWs and EMWs are not close to their critical densities. This condition, however, does not necessarily hold for the dominant wave modes in our theoretical [the peak in figure (5a)] and numerical [figure (4b)] results, where the scattered light of SRS is sufficiently close to its own critical density to violate the WKB approximationMichel (2023). This makes the SRS absolutely unstable and the calculated convective gain from the equation (7) no longer reflects the original meaning of the amplification factor of a delta-function seed. However, we still follow the Rosenbluth gain formula [equation (7)] here. First, except for those dominant modes, this gain formula works well for most matching modes at other densities and can properly reflect their gains. Second, the “erroneous" results of this formula for the dominant absolute SRS modes are actually overwritten by the effects of finite width of the laser beams, as we will explain in the following few paragraphs [equation (9) and figure (5b)].
Using Rosenbluth gain theory, the total gain of the TSI common waves driven simultaneously by TPD and SRS can be evaluated by considering the contributions from both TPD and SRS:
$G_{tot}=G_{S}\times G_{T}$ (8)
where $G_{S}$ and $G_{T}$ are the gains of SRS and TPD, respectively. In 1-dimension cases, they are basically Rosenbluth gains described in equation (7). Please note that achieving $G_{R}$ relies on the complete amplifications process of initial seeds traveling along the density gradient for a finite distance. However, in our simulations using Gaussian beams with finite beam widths, the amplification processes of TPD and SRS seeds which have finite transverse velocities are limited by the transverse size of the laser overlap region [figure (1a)]. Namely, the unstable seeds may leave the laser overlap region and stop growing before reaching $G_{R}$. Therefore, the gain should be truncated to the amplification within the laser overlap region, and the gains $G_{S}$ and $G_{T}$ should be similarly written as
$G_{S}\text{ or }G_{T}=min(G_{R},G_{w})$ (9)
where
$G_{w}=exp[\frac{\gamma_{0}L_{tr}}{max(|V_{1y}|,|V_{2y}|)}]$ (10)
is the gain assuming the unstable seed continues growing by growth rate $\gamma_{0}$ without saturation during propagation until leaving the laser overlap region from the sides. Here, $L_{tr}$ is the transverse size of the laser overlap region, and $V_{1y}$ and $V_{2y}$ respectively represent the group velocities of the two daughter waves along the transverse ($y-$) direction. Please note that that the instability growth rate $\gamma_{0}$ for homogeneous plasmas is also the temporal growth rate of the seed propagating in inhomogeneous plasmasRosenbluth, White, and Liu (1973). Thus, the minimum value of $G_{R}$ and $G_{w}$ should represent the overall gain of the common waves.
Figure 5: The calculated (a) $G_{tot}$, (b) $G_{R}$ and $G_{w}$ of TPD and SRS for simulation parameter set (i) in table 1 with $\theta_{in}=52^{\circ}$. The Rosenbluth gain $G_{R}$ of SRS (blue solid line) is greater than $G_{w}$ (blue dashed line) near the narrow peak, and does not contribute to $G_{tot}$.
With the above analysis, we calculate $G_{tot}$ as a function of $n_{e}$ with $T_{e}=3keV$, $L_{n}=150\mu m$ [table 1(i)] with $\theta_{in}=52^{\circ}$. The results are plotted in figure 5(a). We find that the $G_{tot}$ reaches the highest peak within a narrow density region around $0.217n_{c}$ with $k_{x}\approx 0.54\text{ to }0.57\omega_{0}/c$ and $k_{y}\approx 0.91\omega_{0}/c$. Both the location and the wave number agree very well with the strongest EPW modes observed in figure (4b) and (6a). One may notice another neighboring peak close to the common wave with similar $k_{x}$ but slightly larger $k_{y}\approx 0.98\omega_{0}/c$. This signal is caused by the absolute SRS mode driven by single beams, and is discussed at the end of this section.
Figure 6: The $k_{y}-n_{e}$ spectra for (a) simulation (i) and (b) simulation (v) [table 1] with $\theta_{in}=52^{\circ}$ during 5-6 ps. The vertical dashed lines mark the location of the maximum $G_{tot}$ [figure (5a)].
The formation of the common-wave peak relies on the local maximum point of SRS gain $G_{S}$ , as the TPD gain $G_{T}$ does not show any local maximum point [figure (5b)]. However, although the SRS Rosenbluth gain $G_{R}$ reaches an extremely high peak around the same location due to $\kappa^{\prime}=\frac{d}{dx}(k_{0x}-k_{1x}-k_{2x})\approx 0$, it is not the cause of $G_{S}$ peak, as it is overwritten by the local $G_{w}$, as shown by the blue solid and dashed lines in figure (5b).
Additionally, we clarify that the local maximum $G_{w}$ is not the one with SRS scattered light pointing strictly parallel ($0^{\circ}$) to the y-direction . Instead, the strongest SRS scattered light travels along $\theta_{s}\sim\pm 95^{\circ}$ to the positive x-direction, according to our calculations. This qualitatively agrees with the simulation results in figure 4(b), where the SRS scattered light covers the band region starting from ($k_{x}\sim 0$, $k_{y}\sim 0.21\omega_{0}/c$) and extending to negative $k_{x}$, maintaining the same $k_{y}$ during the propagation. This phenomenon is more clearly depicted in the $k_{y}-n_{e}$ spectra of $E_{y}$ in figure (6), where the horizontal lines with $k_{y}\sim 0.21\omega_{0}/c$ represent the propagation of the scattered light towards the negative x-direction. The extension of these lines towards the left boundary of the simulation domain is not expected for the scattered light with $\theta=90^{\circ}$, as such light would propagate exclusively along the y-direction, with no component in the x-direction.
In addition to the example discussed above, our theoretical predictions of the peak common wave locations align well with the PIC simulation results under various laser plasma conditions. Figure (6) presents the $k_{y}-n_{e}$ phase spaces of $E_{y}$ overlaid with the peak locations of $G_{tot}$ for the simulation (i) and (v) [table 1] with $\theta_{in}=52^{\circ}$. In the two $k_{y}-n_{e}$ figures, the modes near (a) $0.22n_{c}$ and (b) $0.23n_{c}$ with $k_{y}\sim 0.9\omega_{0}/c$ are the TSI common EPWs, while those at the same density but with much larger $k_{y}$ correspond to the other paired TPD modes. These figures show that the $G_{tot}$ peak locations consistently match the theoretical TSI common wave locations.
In figure (6a), we observe additional modes in the lower density region, exhibiting slightly larger $k_{y}$ than the common waves. In the $k_{x}-k_{y}$ phase space, these modes are situated at ($k_{x}\sim 0.54\omega_{0}/c$, $k_{y}\sim 0.98\omega_{0}/c$) [figure 3(d) and 4(b)], near the band signals of the common waves. We identify these modes as EPWs of absolute SRS modesAfeyan and Williams (1985), with corresponding scattered light at ($k_{x}\sim 0$, $k_{y}\sim 0.28\omega_{0}/c$), driven by single laser beams near the incident plane at $n_{e}\geq 0.2n_{c}$. This absolute SRS occurs when $\kappa^{\prime}$ close to zero. For the simulation (i) with $\theta_{in}=52^{\circ}$ at $n_{e}=0.2n_{c}$, our calculations of absolute SRS indicate ($k_{x}\approx 0.53\omega_{0}/c$, $k_{y}\approx 0.97\omega_{0}/c$) for the EPW modes and ($k_{x}\approx 0.02\omega_{0}/c$, $k_{y}\approx 0.27\omega_{0}/c$) for the scattered light, which are in good agreement with the above simulation results. Absolute SRS modes can develop when the laser-plasma conditions surpass the instability thresholdAfeyan and Williams (1985). In our simulations, they occur mildly in some cases for which the threshold is slightly exceeded while not observed in others, suggesting they play a negligible role in both hot electron generation and laser reflection.
III.3 The angular characteristics of hot electrons for different $\theta_{in}$
The driven EPWs can effectively accelerate background thermal electrons with velocity close to the EPW phase velocity $v_{ph}$. Therefore, the hot electron characteristics are closely tied to the EPW spectra. For smaller $\theta_{in}=24^{\circ}$, the EPW spectra exhibit a two-island structure, as shown in figure (3c). The main island, centered around the origin, contains EPWs with a broad range of wave vectors, which complicates the analysis of the resulting hot electrons. Fortunately, most of these modes have $|k|<0.5\omega_{0}/c$, corresponding to $v_{ph}\geq c$, which is significantly greater than the electron thermal velocities, thus inefficient for generating hot electrons. Instead, the smaller island with $k_{x}>1\omega_{0}/c$ and $k_{y}\sim 0$, leading to $v_{ph}\sim 0.5c$, can efficiently capture and accelerate background electrons with temperatures around a few $keV$. Correspondingly, as shown in figure (7), the hot electrons accelerated in the case of $\theta_{in}=24^{\circ}$ exhibit a collimated feature in the angular distribution. Also, the same feature is observed for $\theta_{in}$ =32 when the side TPD and SRS common waves are not significantly driven.
Figure 7: The angular distribution of hot electrons for the simulation (i) in table 1 with different laser incident angle $\theta_{in}$’s.
For larger $\theta_{in}=52^{\circ}$, the EPW spectra display a multi-peak structure (figure 3d and 4b). Thus, the hot electron characteristics should be determined by these peaks. According to our analysis in the previous section, these peaks mainly correspond to the modes of TPD and SRS co-driven by duel beams , and primarily point sideways. Specifically, the common EPW propagates at $\sim 60^{\circ}$ to the x-direction, while the other TPD daughter wave has $k_{x}\sim 0$ and $k_{y}\sim 1.6\omega_{0}/c$, perpendicular to the x-direction. These two EPWs should generate hot electrons with high divergent angles. This feature is reflected in the hot electron angular distributions for $\theta_{in}\geq 48^{\circ}$ [figure (7)] which exhibit pronounced side peaks at $\sim 72^{\circ}$.
Furthermore, the hot electron angular distributions also show a peak at $0^{\circ}$ along the x-direction, corresponding to collimated hot electrons. However, the EPW spectrum does not indicate any EPW modes traveling along the x-direction [figure (4b)]. To address this discrepancy, we find that the collimated hot electrons can be generated by two common EPW modes traveling oblique to x-direction at symmetric angles. Owing to symmetric laser beams, the common EPWs pointing up-right and down-right share the same $k_{cx}$ (the wave number along x-direction) with positive and negative $k_{cy}$ (the wave number along y-direction). The common EPWs along the two directions can form a forward-going phase of electric field, which can capture and accelerate background electrons similar to an EPW traveling along the x-direction. This can be demonstrated by a brief derivation. We assume $E_{cx1}(x,t)$ and $E_{cx2}(x,t)$ represent the electric fields along the x-direction of the two common EPWs, given by:
$E_{cx1}(x,t)=E_{cx0}sin(k_{cx}x+k_{cy}y-\omega_{c}t+\phi_{c1})$ (11)
$E_{cx2}(x,t)=E_{cx0}sin(k_{cx}x-k_{cy}y-\omega_{c}t+\phi_{c2})$ (12)
where $E_{cx0}$ and $\omega_{c}$ are the field amplitude and frequency of the common waves, respectively, with random phases $\phi_{c1}$ and $\phi_{c2}$. The total electric field $E_{cx}$ caused by the common EPWs becomes
$E_{cx}$ $=E_{cx1}(x,t)+E_{cx2}(x,t)$
$=2E_{cx0}sin(k_{cx}x-\omega_{c}t+\phi_{c1}/2+\phi_{c2}/2)$
×$\times cos(k_{cy}y+\phi_{c1}/2-\phi_{c2}/2)$ (13)
which consists of a forward-propagating electric field $sin(k_{cx}x-\omega_{c}t)$ with phase velocity $v_{cph}=\omega_{c}/k_{cx}$, modulated along the y-direction by $cos(k_{cy}y)$. For this field $E_{cx}$, the background electrons can be captured and accelerated as if they are being driven by an EPW with $v_{ph}=v_{cph}$ along the x-direction.
The electric field structure described by the equation (III.3) is observed in the overlap region of the common waves propagating up-right and down-right. This region is situated around $n_{e}\sim 0.217n_{c}$ or $x=193c/\omega_{0}$ and $y=1640c/\omega_{0}\sim L_{y}/2$. In this area, although the fields of common waves are always mixed with the fields from the incident laser beams, their spatial spectra are clearly separated. Therefore, we can isolate the spatial distribution of only the common wave fields by applying spectral filtering in the Fourier-transformed space and then performing an inverse Fourier transform. Figure (8a) presents the filtered common wave fields at $t=1.9ps$, where both longitudinal and transverse modulations are visible around ($k_{x}=0.54\omega_{0}/c$, $k_{y}=0.91\omega_{0}/c$), in agreement with the waveform predicted by equation (III.3). This common wave field should propagate along x-direction and is capable of capturing and accelerating background electrons to collimated hot electrons.
The captured electrons being accelerated by this field can be visualized through the energy density profile of the electrons, defined as:
$\epsilon_{e}(i,j)=\sum_{k=1}^{N_{ij}}w_{e}(\gamma-1)$ (14)
where $\epsilon_{e}$ represents the electron energy density in the cell indexed by $(i,j)$. The number of electrons in the cell is denoted by $N_{ij}$, with $w_{e}$ and $\gamma$ referring to the charge weight and the relativistic gamma factor of the $k$-th electron, respectively. In the absence of the accelerated background electrons, $\epsilon_{e}$ is proportional to $n_{e}T_{e}$, which is approximately $1.9\times 10^{-3}$ at $n_{e}\sim 0.217n_{c}$. When common waves begin accelerating electrons, the local $\epsilon_{e}$ will increase in regions where these electrons are located. Figure (8b) depicts the spatial distribution of $\epsilon_{e}$, where the yellow bands with $\epsilon_{e}>1.9\times 10^{-3}$ corresponds to the areas where accelerated electrons are located.
By comparing the common wave fields to the locations of the accelerated electrons shown in figure (8a) and (8b), it becomes evident that the hot electrons are concentrated in bands positioned between the peaks of the common wave fields. The parallel yellow bands along the y-direction suggest that these electrons are being accelerated by the wave fields and are moving along the x-direction. These findings indicate that the symmetric, divergent common waves can produce collimated hot electrons.
Figure 8: (a) The $E_{cx}$ fields [equation (III.3)] of two common waves and electron energy density $\epsilon_{e}$ [equation (14)] in the overlap region near $n_{e}=0.217n_{c}$ at $t=1.9ps$ for the PIC simulation (i) [Table 1] with $\theta_{in}=52^{\circ}$. The data in (a) are obtained by filtering the space spectra of the common waves out of the total electric field using Fourier and inverse-Fourier transforms.
III.4 The energy scaling of hot electrons with different divergent angles
As we discussed in the previous section, hot electrons with different divergent angles are generated through varied mechanisms. Thus, they may follow different energy scalings. According to the divergent angle $\theta_{e}$, we classify the hot electrons into three categories: collimated ($\theta_{e}<30^{\circ}$), intermediate ($30^{\circ}<\theta_{e}<45^{\circ}$) and divergent ($\theta_{e}>45^{\circ}$). Figure (9) plots the energy fractions of these groups against $G_{tot}$ for the simulations listed in table 1, specifically for $\theta_{in}\geq 40^{\circ}$. Smaller $\theta_{in}$ values correspond to the dominant forward dual-TPD common wave mechanism [section III.2], rather than the side TPD-SRS common waves, which $G_{tot}$ is based upon. Among these simulations, the results of simulation (ii) and (iv) with $\theta_{in}\geq 40^{\circ}$ are not given, as the $f_{hot}$ has not reached quasi-steady state till the end of the simulation time due to the weak instability growth.
Our findings show that both the collimated and divergent hot electrons tend to have higher energy than the intermediate group, consistent with the multi-peak angular distributions in figure (7). In this log-log plot, the $f_{hot}$ values of all three groups increase quasi-linearly with $G_{tot}$, but with notably different slopes. This behavior suggests the following approximate scaling:
$f_{hot}\approx CG_{tot}^{\alpha}$ (15)
where $\alpha=$ 0.19, 0.28 and 0.49 , and the coefficient $C=$ $1.7\times 10^{-2}$, $4.8\times 10^{-3}$ and $8.5\times 10^{-3}$ for the collimated, intermediate and divergent groups of hot electrons, respectively.
The different $\alpha$ values indicate that $f_{hot}$ exhibits varying sensitivity to $G_{tot}$, which could be related to the differences in acceleration efficiency among the groups. Divergent electrons are accelerated by the common wave and the paired side TPD mode. In the example of the simulation (i) with $\theta_{in}=52^{\circ}$, the two modes have $|k|\sim 1.1\omega_{0}/c$ and $1.6\omega_{0}/c$, corresponding to phase velocities of $0.45c\sim 5.9v_{te}$ and $0.31c\sim 4.0v_{te}$, respectively. In contrast, according to the equation (III.3), the collimated hot electrons are accelerated by the moving electric field with phase velocity of $v_{cph}=\omega_{c}/k_{cx}\sim 0.88c\sim 11.5v_{te}$, much faster than that for divergent hot electrons. These results show that divergent electrons are more easily accelerated due to the lower phase velocity of EPW modes compared to collimated electrons. This difference may account for the varying sensitivity of $f_{hot}$ in response to changes in the amplitude of EPW modes, which is directly related to $G_{tot}$. A detailed study on the reasons behind these different $\alpha$ values will be the focus of future research.
Figure 9: Scaling relations between energy of hot electrons ($f_{hot}$ normalized by incident beams) with (a) $\theta_{e}<30^{\circ}$, (b) $30^{\circ}<\theta_{e}<45^{\circ}$ and (c) $\theta_{e}>45^{\circ}$) and $G_{tot}$ of common wave driven by TPD and SRS.
IV Considering Realistic Experimental Conditions
IV.1 Effects of asymmetric incidence
Figure 10: (a) The space spectra of $E_{y}$ for the simulation (b) [table 2] during the simulation time of 5-6ps. (b) The angular distributions of hot electrons for all the simulations in table 2 with different $\theta_{in}$’s.
In realistic experiments on large-scale laser facilities, despite the symmetric configuration of the beams, the orientation of density contours evolves in space and time during the target surface ablation. As a result, it is common for laser beams to deviate from symmetry relative to the normal of the density contours near $n_{c}/4$. Therefore, hot electron generation under asymmetric incidence is crucial for accurately assessing $f_{hot}$ in experimental evaluations. In such scenarios, the matching ellipses [Figure (4a)] for both TPD and SRS mechanisms continue to define the common wave modes responsible for the TSI mechanism. The cumulative wave amplification at various densities can still be scaled by the total gain, $G_{tot}$, as derived from equations (7) to (10).
We calculate two new cases using the same physical parameters as case (i) in table (1) but with the laser incident angles directed up-right and down-right ($\theta_{in,up}$, $\theta_{in,down}$) changed from symmetric ($52^{\circ}$, $-52^{\circ}$) to asymmetric configurations of ($48^{\circ}$, $-56^{\circ}$) and ($44^{\circ}$, $-60^{\circ}$). These shifts move the beams away from the density gradient direction by $4^{\circ}$ and $8^{\circ}$, as detailed in table (2).
Table 2: The incident angles and calculated $G_{tot,up}$ and $G_{tot,down}$ for asymmetric incident laser beams.
Index $\theta_{in,up}(^{\circ})$ $\theta_{in,down}(^{\circ})$ $G_{tot,up}$ $G_{tot,down}$
(a) 52 -52 25 25
(b) 48 -56 28 23
(c) 44 -60 32 20
The computed total gains, $G_{tot}$, for the common waves propagating up-right and down-right, denoted as $G_{tot,up}$ and $G_{tot,down}$, are presented in table (2) for all the three combinations of ($\theta_{in,up}$, $\theta_{in,down}$). From these results, we observe that the asymmetric incident beams lead to variations of up to $28\%$ in $G_{tot,up}$ and $G_{tot,down}$ when compared to the symmetric case. These variations results in differing amplitudes for the two common wave modes and their associated TPD modes, leading to asymmetric angular distributions of hot electrons. As illustrated in Figure (10a), the amplitude of the common wave propagating up-right (black arrow) is noticeably stronger than that of the wave propagating down-right. For the stronger up-right common wave, the corresponding SRS-scattered light (orange arrow) and the paired TPD mode (white solid arrow), both of which point near the negative $k_{y}$ direction, exhibit significantly greater strength than those associated with the other down-right common wave.
Due to the asymmetry in $G_{tot,up}$ and $G_{tot,down}$, the angular distributions of hot electrons exhibit an asymmetric shape, as illustrated in figure (10b)]. Interestingly, although $G_{tot,up}>G_{tot,down}$ [table (2)] suggests a stronger common wave propagating in the up-right direction (positive angles), a greater number of hot electrons are produced along negative angles [figure (10b)]. This apparent contradiction arises because the hot electrons are generated by both the common wave and the paired TPD modes. The two waves travel oppositely in y-direction, meaning that hot electrons with positive angles are influenced by the stronger common wave propagating up-right and the weaker TPD mode paired with the down-right common wave, and vice versa. As a result, it is not immediately clear whether more hot electrons will be generated at positive or negative angles. However, the results indicate a greater concentration of hot electrons at negative angles, implying that hot electron generation is more sensitive to the TPD modes with higher $|k|$ and lower phase velocity.
Moreover, the total hot electron energy increases by approximately 20% for asymmetric $\theta_{in}$’s, as the surplus of hot electrons on one side outweighs the reduction on the other.
IV.2 Experimental observability of the TSI mechanism
Since the scattered light in the TSI mechanism is essentially enhanced SRS scattered light, it should be detectable in experiments. For the simulation (i) [table 1] with symmetric incident angle $\theta_{in}=52^{\circ}$, if we assume $n_{e}$ only varies along $x-$ direction in space, the scattered light with a frequency of $\omega\sim 0.512\omega_{0}$ and wave vector component $k_{y}\sim 0.212\omega_{0}/c$ will ultimately have a scattering angle of only $\sim 24^{\circ}$ relative to the negative x-direction after exiting into the vacuum. This scattered light has the same frequency as the collective Thomson scattering of the incident light on the common EPW mode. Once diagnosed, it should mix with the double-band structure of the TPD $\omega_{0}/2$ spectra, and strengthen the blue-shift component. This qualitatively agrees with the observations of significantly stronger blue-shift components compared to red-shift components off the target normal direction in previous OMEGA experimentsSeka et al. (2014).
IV.3 The actual incident angles in the vacuum
It is important to clarify that all $\theta_{in}$ values in this article refer to the incident angles at the density $n_{e}\sim 0.2n_{c}$ on the incident plane of our PIC simulations. These angles typically differ from the actual vacuum incident angles, $\theta_{in,vac}$, which are the true parameters defining the beam geometry. We now discuss the relationship between $\theta_{in,vac}$ and $\theta_{in}$ by analyzing two extreme cases.
In the first case, where the plasma density is spherically symmetric with the incident beams directed toward the center, we have $\theta_{in,vac}\sim\theta_{in}$. In contrast, for a 1D planar geometry where $n_{e}$ varies only along the x-axis, the relationship between $\theta_{in,vac}$ and $\theta_{in}$ can be determined by the fact that $k_{y}$ is constant and by using the light wave dispersion relation at $n_{e}=0.2n_{c}$:
$\theta_{in,vac}\approx arcsin[0.89sin\theta_{in}]$ (16)
This yields pairs of ($\theta_{in,vac}$, $\theta_{in}$) such as ($50^{\circ}$, $60^{\circ}$), ($43^{\circ}$, $50^{\circ}$), and ($35^{\circ}$, $40^{\circ}$). Generally, for a given $\theta_{in,vac}$, the effective incident angle for TPD should lie between $\theta_{in,vac}$ and the calculated $\theta_{in}$ from equation (16).
IV.4 The connection between picosecond PIC simulation results to nanosecond hydrodynamic evolution
A key question is the extent to which our PIC simulation results, obtained on a picosecond time scale, can inform the realistic implosion process occurring over several nanoseconds. The significant disparity between these time scales naturally raises concerns about the applicability of our results to real-world experiments. To address this, it is important to note that as long as a quasi-steady state is achieved during the short-time-scale process, quantities from this state can reasonably be coupled to the long-time-scale process with slowly evolving hydrodynamic parameters. This is because the long-time-scale process is not influenced by transient growth or decay occurring before the system reaches its quasi-steady state. Following this principle, previous studiesFollett et al. (2017) have successfully predicted the time evolution of hot electron energy in nanosecond-scale experiments by using results from many hybrid fluid-particle simulations (each runs for 20 ps) that model hot electron generation under various laser-plasma conditions throughout the entire laser pulse and focal spot. Another successful example is the development of iterative methods that couple 15 ps-PIC and 2 ns-hydro simulations to assess the impact of TPD-induced strong pump depletion due to anomalous laser absorption and hot electron generation on hydrodynamic evolutionLi, Hu, and Ren (2017).
In the context of our simulations, since the results clearly exhibit quasi-steady states [figure (2)], we expect our findings to reliably reflect the hot electron characteristics for the corresponding physical parameters and beam geometry of the laser pulses and plasmas.
V conclusion
In this study, we investigate the angular variations of hot electrons generated by multiple laser beams with varying incident angles. We identify a novel multi-beam collaborative process involving two-plasmon decay (TPD) and stimulated Raman scattering (SRS), termed TSI, which produces hot electrons with distinct divergent and collimated components.
We demonstrate that the key factor influencing the acceleration of these components is the common Langmuir wave associated with TSI. By properly modeling the common wave gain, we obtain scaling relations between the gains and the energy fractions of hot electrons at different angles. Our results show that the energy of divergent hot electrons is more sensitive to variations in the common wave gain. Furthermore, when the density gradient deviates from the bisector of the laser beams— a common occurrence in large-scale laser facilities—our study reveals that asymmetric beam incidence results in unequal common wave gains, leading to asymmetric angular distributions of hot electrons. The stronger peak is observed in the direction opposite to the laser beam with the larger incident angle. These findings highlight potential preheating concerns due to divergent and collimated hot electrons in laser-plasma configurations with high gains in future DCI experiments.
Finally, we observe that the scattered light associated with the TSI mechanism qualitatively agrees with previous experimental findings. This scattered light shares the frequency of collective Thomson scattered light from the common wave. Although the onsite scattering angle is large (approximately $95^{\circ}$), the direction of the scattered light shifts towards the density gradient when propagating towards vacuum, with final angle of about $24^{\circ}$ in our simulations for 1D plasma density variations. This scattered light merges with the double-band structure of the TPD $\omega_{0}/2$ spectrum, enhancing the blue-shift component, consistent with prior OMEGA experimentsSeka et al. (2014) that observed significantly stronger blue-shift than red-shift signals when viewed off the target normal.
Acknowledgements.
This research was supported by the National Natural Science Foundation of China (NSFC) (Grant Nos. 12275269, 12175229, 12275032), by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Nos. XDA25010200 and XDA25050400). The numerical simulations in this paper were conducted on Hefei advanced computing center.